# A Learning-Based Framework for Safe Human-Robot Collaboration with Multiple Backup Control Barrier Functions

### v1

@@ Abstract
Ensuring robot safety in complex environments is a difficult task due to actuation limits, such as torque bounds. This paper presents a safety-critical control framework that leverages learning-based switching between multiple backup controllers to formally guarantee safety under bounded control inputs while satisfying driver intention. By leveraging backup controllers designed to uphold safety and input constraints, backup control barrier functions (BCBFs) construct implicitly defined control invariance sets via a feasible quadratic program (QP). However, BCBF performance largely depends on the design and conservativeness of the chosen backup controller, especially in our setting of human-driven vehicles in complex, e.g, off-road, conditions. While conservativeness can be reduced by using multiple backup controllers, determining when to switch is an open problem. Consequently, we develop a broadcast scheme that estimates driver intention and integrates BCBFs with multiple backup strategies for human-robot interaction. An LSTM classifier uses data inputs from the robot, human, and safety algorithms to continually choose a backup controller in real-time. We demonstrate our method’s efficacy on a dual-track robot in obstacle avoidance scenarios. Our framework guarantees robot safety while adhering to driver intention.
@@ I Introduction
Consequently, the CBF framework has been extended to include actuation capability, such as torque limits, through the use of backup control barrier functions (BCBFs) [12, 13, 14]. BCBFs rely on the formulation of a backup controller, which is designed with input limits in mind to guarantee safety. The backup controller typically involves a simple saving maneuver, such as coming to a stop, turning at a maximum rate, or hovering. By calculating the future backup trajectory of the system, one can analyze future safety of the robot and incorporate this information into optimization-based controllers like quadratic programs (QPs). Consequently, when constructing safety filters using the BCBF method, the system’s conservativeness is a strong function of the control engineer’s choice of backup policy.
@@
To address this limitation, recent work examines the use of multiple backup strategies in the BCBF framework, since multiple strategies can help overcome the conservativeness of a single strategy. In [15], an algorithm is used to evaluate the BCBF method with multiple backup controllers, such that the one with the least control intervention can be chosen. However, with many backup controllers, this method could be computationally infeasible. In [16] and [17], different maneuvers are proposed to increase the reachability of a given backup controller, where a switching algorithm chooses a different maneuver if it is no longer possible to perform the current maneuver. While this method used multiple backup maneuvers with more computationally efficient ways of solving BCBFs, it is better suited to fully autonomous systems. It does not explicitly provide, in a human-robot interaction context, a way to take into account the intentions of a human vehicle driver or human companion. Nor does it account for the fact that the human operator may have better situational awareness than the autonomous controller. Even though autonomous selection may be suboptimal at times, the driver does not have the authority to bypass the chosen backup control strategy. Additionally, it cannot know if the driver has suddenly changed his or her goals. Generally, the types of safe behavior utilized by the BCBF cannot be tuned.
@@
There has been impressive previous work to include human preferences in safety filters. For instance in [18], preference-based learning is used to adapt tuneable parameters of a CBF to user preferences for autonomous obstacle avoidance on a quadruped. However, input bounds are not formally considered, which could potentially result in infeasiblity and consequently unsafe trajectories. While there has also been work on preference based reinforcement learning (pbRL) [19], these methods must sample trajectories within the expected environment [20]. Since the drivers of teleoperated systems must provide control inputs at each timestep, trajectory sampling becomes difficult to complete due to the human-in-the-loop nature of the problem.
@@
In order to intelligently pick backup controllers, we propose a system that utilizes a long short term memory (LSTM) and a deep neural network (DNN) component to learn a reward corresponding to each backup controller. LSTMs have been used in real-time driver intention and maneuver tracking, and have shown to be successful at regression and clasification tasks with trajectory input [21, 22, 23]. A simple switching law is derived by using the controller with the largest reward in the BCBF framework. This system learns to estimate driver intention by training on example trajectories, labeled with the correct choice of backup controller, as provided by the driver. In our work, we show that
@@
    The LSTM-DNN switching architecture effectively learns the correct backup controller.
@@
This paper is organized as follows. Section II presents preliminaries on CBFs, BCBFs, DNNs and LSTMs. Section III presents our contributions, including a description of system architecture. We present the implementation of our framework on hardware in Section IV and discuss experimental results in Section 5. Section VI concludes the paper.
@@ II-B Backup Control Barrier Functions
BCBFs are motivated by the fact that verifying (5), which is required for the feasibility of (6), may be challenging for a particular choice of ${h}$, especially with bounded inputs. Consider input bounds with component-wise hard constraints:
@@
where ${u_{\rm max}\in\mathbb{R}^{m}_{>0}}$. The backup set method [12, 13, 14] addresses this feasibility problem by designing implicit control invariant sets and safe controllers through a CBF framework.
We consider a backup set ${\mathcal{C}_{\rm b}\!\subset\!\mathbb{R}^{n}}$ defined as the 0-superlevel set of a continuously differentiable function ${h_{\rm b}:\mathbb{R}^{n}\!\to\!\mathbb{R}}$:
@@
Designing a control invariant backup set ${\mathcal{C}_{\rm b}}$ is generally easier than verifying if ${\mathcal{C}}$ is control invariant. However, the methods used to develop ${\mathcal{C}_{\rm b}}$ may result in a conservative set [5]. We can reduce conservatism by expanding the backup set. To achieve this, we use the backup trajectory ${\phi_{\rm b}(\tau,x)}$ over a finite time period ${\tau\in[0,T]}$ with some ${T\in\mathbb{R}_{>0}}$. Note that ${{\phi_{\rm b}}(\tau,x)}$ is the flow map of the system under the backup controller with the initial condition ${x}$. Then we define a larger control invariant set, called $\mathcal{C}_{E}$, such that ${\mathcal{C}_{\rm b}\subseteq\mathcal{C}_{E}\subseteq\mathcal{C}}$:
$\begin{split}\mathcal{C}_{E}\triangleq\left\{x\in\mathcal{C}~{}\Bigg{|}\begin{array}[]{ll}\overbrace{h(\phi_{\rm b}(\tau,x)}^{\triangleq\bar{h}_{\tau}(x)}\geq 0,~{}\forall\tau\in[0,T],\\ \underbrace{h_{\rm b}(\phi_{\rm b}(T,x))}_{\triangleq\bar{h}_{\rm b}(x)}\geq 0\end{array}\right\}.\end{split}$ (11)
@@
[…]
@@ II-C Deep Neural Networks
where $\upsilon_{\jmath}$ is the output of the $\jmath$-th layer of the DNN, and ${\upsilon_{0}=\mathbf{\bar{x}}}$ is the input to the DNN and ${\upsilon_{N_{L}}=\mathbf{\bar{y}}}$ is the output of the DNN, respectively. Each hidden layer has ${n_{\jmath}\in\mathbb{N}}$ hidden neurons. ${\mathbf{\bar{W}}_{\jmath}\in\mathbb{R}^{n_{\jmath}\times n_{\jmath-1}}}$ and ${b_{\jmath}\in\mathbb{R}^{n_{\jmath}}}$ are weight matrices and bias vectors for the $\jmath^{{th}}$ layer, respectively. ${\pi_{\jmath}\triangleq\left[\varrho_{\jmath},\cdots,\varrho_{\jmath}\right]}$ is the concentrated activation functions of the $\jmath^{{th}}$ layer wherein ${\varrho_{\jmath}:\mathbb{R}\rightarrow\mathbb{R}}$ is the activation function, such as sigmoid or ReLU:
@@
Specifically, we use ${\varrho_{\textrm{sigmoid}}}$ function in our experiments.
II-D Long Short Term Memories
Long Short Term Memories (LSTMs) are a class of Recurrent Neural Networks (RNNs) which have several advantages over DNNs when learning temporally related data. LSTMs also remedy the exploding/vanishing gradient problem of conventonal RNNs by providing a recurrent path for long term memories and another path for short term memories, as the name suggests. LSTM networks [24] learn to map input ${\mathbf{\tilde{x}}=\left[\mathbf{\tilde{x}}_{n\tilde{T}}\right]\in\mathbb{R}^{\tilde{T}n_{x}}}$ to output sequence ${\mathbf{\tilde{y}}=\left[\mathbf{\tilde{y}}_{n\tilde{T}}\right]\in\mathbb{R}^{\tilde{T}n_{y}}}$ for ${n=0,1,\ldots,N}$, where ${\tilde{T}\in\mathbb{R}_{>0}}$ is the sampling time of the LSTM. To learn temporal relationships in this mapping, LSTM units contain a memory cell and four gates: the forget $\mathbf{\tilde{f}}$, input $\mathbf{\tilde{i}}$, output $\mathbf{\tilde{o}}$, and cell $\mathbf{\tilde{c}}$ gates. The recursively defined equations give the outputs of the gate functions:
$\begin{split}\mathbf{\tilde{f}}_{n}&=\varrho_{\tilde{f}}(\mathbf{\tilde{W}}_{\mathbf{\tilde{f}}}\mathbf{\tilde{x}}_{n}+\mathbf{\tilde{R}}_{\mathbf{\tilde{f}}}\mathbf{\tilde{h}}_{n-1}+\mathbf{\tilde{b}}_{\mathbf{\tilde{f}}}),\\ \mathbf{\tilde{i}}_{n}&=\varrho_{\tilde{f}}(\mathbf{\tilde{W}}_{\mathbf{\tilde{i}}}\mathbf{\tilde{x}}_{n}+\mathbf{\tilde{R}}_{\mathbf{\tilde{i}}}\mathbf{\tilde{h}}_{n-1}+\mathbf{\tilde{b}}_{\mathbf{\tilde{i}}}),\\ \mathbf{\tilde{o}}_{n}&=\varrho_{\tilde{o}}(\mathbf{\tilde{W}}_{\mathbf{\tilde{o}}}\mathbf{\tilde{x}}_{n}+\mathbf{\tilde{R}}_{\mathbf{\tilde{o}}}\mathbf{\tilde{h}}_{n-1}+\mathbf{\tilde{b}}_{\mathbf{\tilde{o}}}),\\ \mathbf{\tilde{g}}_{n}&=\varrho_{\tilde{g}}(\mathbf{\tilde{W}}_{\mathbf{\tilde{g}}}\mathbf{\tilde{x}}_{n}+\mathbf{\tilde{R}}_{\mathbf{\tilde{g}}}\mathbf{\tilde{h}}_{n-1}+\mathbf{\tilde{b}}_{\mathbf{\tilde{g}}}),\end{split}$
where $n$ and $n-1$ refer to current and previous time steps, ${\varrho_{\tilde{f}},\varrho_{\tilde{i}},\varrho_{\tilde{o}}}$ are the activation functions for input, output and forget gates, and are generally chosen as ${\varrho_{\textrm{sigmoid}}}$. ${\varrho_{\tilde{g}}}$ is the activation function for the memory cell state, which is $tanh$ for general cases, ${\mathbf{\tilde{x}}_{n}}$ is the normalized input, ${\mathbf{\tilde{h}}_{n-1}}$ is the hidden state, and ${\mathbf{\tilde{W}}}$, ${\mathbf{\tilde{R}}}$, ${\mathbf{\tilde{b}}}$ are the weight matrices for $\mathbf{\tilde{i}}$ and the recurrent weight matrices, and bias vectors. The update equations for ${\mathbf{\tilde{h}}_{n-1}}$ and the cell state $\mathbf{\tilde{c}}_{n}$ given by
$\mathbf{\tilde{h}}_{n}=\mathbf{\tilde{o}}_{n}\odot\varrho_{\tilde{g}}(\mathbf{\tilde{c}}_{n});\ \ \mathbf{\tilde{c}}_{n}=\mathbf{\tilde{f}}_{n}\odot\mathbf{\tilde{c}}_{n-1}+\mathbf{\tilde{i}}_{n}\odot\mathbf{\tilde{g}}_{n},$
where $\odot$ is the Hadamard (element-wise) product.
Then, with a fully connected layer, weight matrix ${\mathbf{\tilde{W}}_{y}}$ and bias vector ${\mathbf{\tilde{b}}_{y}}$, the output of the LSTM network is given by
$\mathbf{\tilde{y}}=\mathbf{\tilde{W}}_{y}\mathbf{\tilde{h}}_{n}+\mathbf{\tilde{b}}_{y}.$
Notably, ${\mathbf{\tilde{W}}}$ and ${\mathbf{\tilde{b}}}$ do not vary with $n$, as the feedback loop nature of LSTMs and RNNs allow the same weights and biases to be used for each timestep in a sequence of data. This allows LSTMs and RNNs to be applied on input sequences of varying lengths without changes in architecture.
III Intention Estimation
To improve the reachability of multiple BCBFs through intention estimation, a deep LSTM-DNN network translates observational data from the driver, safety systems, and robot state into the choice of $\mathbf{k_{b}}$.
@@ III Intention Estimation
  • $\gamma_{t}\in\mathbb{R}^{k}$ is a vector of ${k\in\mathbb{N}}$ features: robot state, environment, driver input, and safety data. We maintain an ${H+1}$-length data history ${\vec{\gamma}(t,H)=(\gamma_{t},\gamma_{t-1},\ldots,\gamma_{t-H})}$, Additional details on $\vec{\gamma}(t,H)$ can be found below.
  • $R:\mathbb{R}^{k\times(H+1)}\to\mathbb{R}^{m_{k}}$ is a reward function that maps input history $(\gamma_{t},\gamma_{t-1},\ldots,\gamma_{t-H})$ to a list of $m_{k}$ rewards with values ranging from 0 to 1. Each component of this list corresponds to the integer $i$ of each controller $\mathbf{k_{bi}}\in\mathcal{K}$.
@@
where ${p=\begin{bmatrix}{x}_{I}&{y}_{I}\end{bmatrix}^{\top}}$ is the vehicle’s planar position w.r.t. the inertial frame ${I}$, $\theta$ is vehicle’s yaw angle, ${-v_{\rm max}\leq v\leq v_{\rm max}}$ is its linear velocity and ${-\omega_{\rm max}\leq\omega\leq\omega_{\rm max}}$ is the angular velocity. While we describe our intention estimator in the context of the unicycle model, we believe that our system can generalize to more complex dynamical models. This is the subject of future experimentation as detailed in section VI.
@@
We employ an intention estimation framework to learn the reward function, $R$. Learning is useful in this context since constructing a reward function based on multi-modal input in high dimensions may be challenging. The feature vector at a single timestep, ${\gamma_{t}}$, consists of the robot state and vertical position ${[x^{I},z^{I}]\in\mathbb{R}^{4}}$, robot state and vertical velocities ${[\dot{x}^{I},\dot{z}^{I}]\in\mathbb{R}^{4}}$, driver velocity commands ${u\in U}$, and safety information from evaluating ${h:\mathbb{R}^{3}\to\mathbb{R}}$ at $x$ and at ${x_{g}\in\mathbb{R}^{3}}$, which is an intermediate goal determined by forward integrating the driver’s desired $u$ over a short horizon. Note that $\gamma_{t}$ can contain many more features than detailed here in other implementations, especially if they relate to the safety of the robot. However, the aforementioned features generalize to any robot using our implementation. We wish to learn $R$ by mapping history $\vec{\gamma}(t,H)$ to rewards corresponding to each backup controller $\mathbf{k_{bi}}$ at time $t$. This process allows us to derive switching laws for the chosen backup controller as a function of time. For instance, we may choose the backup controller $\mathbf{k_{bi}}$ with the largest reward at time $t$ where $i=\operatorname*{arg\,max}(R(\vec{\gamma}(t,H))$
@@
We use a deep LSTM network with a DNN decoder to construct $R$ from history $\vec{\gamma}(t,H)$. Deep LSTMs can learn complex temporal relationships by using multiple layers and cyclic connections to understand sequential data, and the DNN decoder maps the output of the LSTM to rewards for each backup controller through several hidden layers. Both the LSTM and DNN utilize dropout for regularization, and the DNN uses ReLU activations in the hidden layers. Furthermore, we use the sigmoidal activation function in the DNN final layer. Thus, the outputted rewards for each backup controller, $\mathbf{k_{bi}}$ at time $t$, can be interpreted very loosely as a likelihood that $\mathbf{k_{bi}}$ is a correct backup controller choice at time $t$. For example, if our framework is highly certain that $\mathbf{k_{b0}}$ is the correct choice of backup controller at time $t$, then the first component of the output of ${R(\vec{\gamma}(t,H))}$ is expected to be close to 1.
@@
Our strategy and architecture enables easy reward engineering for the training dataset. We construct a multi-class dataset by gathering data of the robot driving in suitable environments under the correct backup controllers. During data collection the robot should be operating with a BCBF safety filter as we plan to evaluate the network in safety-critical contexts. During training, the switches between backup controllers are labeled in the dataset by the driver using configurable buttons on the robot ground station. Labeling is completed by assigning $1$ to the driver’s choice of backup controller and 00 to all others for that instance. Since we utilize the sigmoidal activation function in this multi-class classification learning task, we use the softmax loss function to train the network. Indeed, this training procedure makes an implicit assumption that there exists a single, correct choice of backup controller at any given time. While this may not be true in certain situations, we guarantee safe switching as discussed in Section IV-B. Consequently, any choice of correct backup controller, or switch between them, does not void safety guarantees. We discuss discontinuities between switching controllers in section 5, however in short, switching between backup controllers is, in practice, smooth.
To improve training, we utilize the Adam optimizer [25] with a stepped learning rate scheduler that reduces the learning rate every constant step of training epochs. In order to compensate for the time it takes to solve the BCBF-QP, we employ a forecast of roughly double the time it takes our system to solve the BCBF-QP. In other words, before training we shift backup controller labels backwards in time so that the network learns to pre-emptively select the correct backup controller, allowing time for the BCBF-QP to have computed safe control outputs for the new controller when they are needed. We show results from a training episode done on hardware-collected data in Fig. 2, and present our model parameters here. We used a 2 layer LSTM with 100 hidden units with the 12 total input features detailed earlier. Our DNN is composed of two hidden layers of sizes 50 and 25 neurons, respectively. We use a dropout value of 0.2 for the hidden layers of the DNN and a dropout of 0.1 for the LSTM. As indicated in Fig. 2, we achieved $97\%$ accuracy over a dataset of 19000 datapoints collected on hardware. Note that the number of datapoints is not the number of training samples since the LSTM-DNN model learns on sequences. Sequence length was chosen to be 15 timesteps, which corresponds to roughly 0.75 seconds of data, so our dataset contained roughly 1266 trajectories.
III-A Human Interface

### v3

@@ Abstract
Ensuring robot safety in complex environments is a difficult task due to actuation limits, such as torque bounds. This paper presents a safety-critical control framework that leverages learning-based switching between multiple backup controllers to formally guarantee safety under bounded control inputs while satisfying driver intention. By leveraging backup controllers designed to uphold safety and input constraints, backup control barrier functions (BCBFs) construct implicitly defined control invariant sets via a feasible quadratic program (QP). However, BCBF performance largely depends on the design and conservativeness of the chosen backup controller, especially in our setting of human-driven vehicles in complex, e.g, off-road, conditions. While conservativeness can be reduced by using multiple backup controllers, determining when to switch is an open problem. Consequently, we develop a broadcast scheme that estimates driver intention and integrates BCBFs with multiple backup strategies for human-robot interaction. An LSTM classifier uses data inputs from the robot, human, and safety algorithms to continually choose a backup controller in real-time. We demonstrate our method’s efficacy on a dual-track robot in obstacle avoidance scenarios. Our framework guarantees robot safety while adhering to driver intention.
@@ I Introduction
Consequently, the CBF framework has been extended to include actuation capability, such as torque limits, through the use of backup control barrier functions (BCBFs) [12, 13, 14]. BCBFs rely on the formulation of a backup controller, which is designed with input limits in mind to guarantee safety.
The backup controller typically involves a simple saving maneuver, such as coming to a stop, turning at a maximum rate, or hovering. By calculating the future backup trajectory of the system, one can analyze future safety of the robot and incorporate this information into optimization-based controllers like quadratic programs (QPs). Consequently, when constructing safety filters using the BCBF method, the system’s conservativeness is a strong function of the control engineer’s choice of backup policy.
@@
To address this limitation, recent work examines the use of multiple backup strategies in the BCBF framework, since multiple strategies can help overcome the conservativeness of a single strategy. In [15], an algorithm is used to evaluate the BCBF method with multiple backup controllers, such that the one with the least control intervention can be chosen. However, with many backup controllers, this method could be computationally infeasible. In [16] and [17], different maneuvers are proposed to increase the reachability of a given backup controller, where a switching algorithm chooses a different maneuver if it is no longer possible to perform the current maneuver. While this method used multiple backup maneuvers with more computationally efficient ways of solving BCBFs, it is better suited to fully autonomous systems. It does not explicitly provide, in a human-robot interaction context, a way to take into account the intentions of a human vehicle driver or human companion. Nor does it account for the fact that the human operator may have better situational awareness than the autonomous controller. Even though autonomous selection may be suboptimal at times, the backup control strategy is often fixed and independent of the driver’s preferred nature of safety-critical behavior. Generally, the types of safe behavior utilized by the BCBF cannot be tuned, leading to situations where potentially preferred safe behaviors are filtered out by the BCBF framework.
@@
There has been impressive previous work to include human preferences in safety filters. For instance in [18], preference-based learning is used to adapt tuneable parameters of a CBF to user preferences for autonomous obstacle avoidance on a quadruped. However, input bounds are not formally considered, which could potentially result in infeasibility and consequently unsafe trajectories. While there has also been work on preference based reinforcement learning (pbRL) [19], these methods must sample trajectories within the expected environment [20]. Since the drivers of teleoperated systems must provide control inputs at each timestep, trajectory sampling becomes difficult to complete due to the human-in-the-loop nature of the problem.
@@
In order to intelligently pick backup controllers, we propose a system that utilizes a long short term memory (LSTM) and a deep neural network (DNN) component to learn a reward corresponding to each backup controller. LSTMs have been used in real-time driver intention and maneuver tracking, and have shown to be successful at regression and classification tasks with trajectory input [21, 22, 23]. A simple switching law is derived by using the controller with the largest reward in the BCBF framework. This system learns to estimate driver intention by training on example trajectories, labeled with the correct choice of backup controller, as provided by the driver. In our work, we show that
@@
    The LSTM-DNN architecture effectively learns a switching law for the BCBF framework with multiple backup controllers–choosing the controller closest to the driver’s preference.
@@
This paper is organized as follows. Section II presents preliminaries on CBFs, BCBFs, and DNNs. Section III presents our contributions, including a description of system architecture. We present the implementation of our framework on hardware in Section IV and discuss experimental results in Section V. Section VI concludes the paper.
@@ II-B Backup Control Barrier Functions
BCBFs are motivated by the fact that finding a function $h$ satisfying the CBF condition (5), which is required for the feasibility of (6), may be challenging for a particular choice of ${h}$, especially with bounded inputs. Consider input bounds with component-wise hard constraints:
@@
where ${u_{\rm max}\in\mathbb{R}^{m}_{>0}}$.
The backup set method [12, 13, 14] addresses this feasibility problem by designing implicit control invariant sets and safe controllers through a CBF framework.
We consider a backup set ${\mathcal{C}_{\rm b}\!\subset\!\mathbb{R}^{n}}$ defined as the 0-superlevel set of a smooth ${h_{\rm b}:\mathbb{R}^{n}\!\to\!\mathbb{R}}$:
@@
Designing a control invariant backup set ${\mathcal{C}_{\rm b}}$ is generally easier than verifying if ${\mathcal{C}}$ is control invariant. However, the methods used to develop ${\mathcal{C}_{\rm b}}$ may result in a conservative set [5]. We can reduce conservatism by expanding the backup set. To achieve this, we use the backup trajectory ${\phi_{\rm b}(\tau,x)}$ over a finite time period ${\tau\in[0,T]}$ with some ${T\in\mathbb{R}_{>0}}$.
Note that ${{\phi_{\rm b}}(\tau,x)}$ is the flow map of the system under the backup controller with the initial condition ${x}$. We define a larger control invariant set, called $\mathcal{C}_{E}$, such that ${\mathcal{C}_{\rm b}\!\subseteq\!\mathcal{C}_{E}\!\subseteq\!\mathcal{C}}$:
$\begin{split}\mathcal{C}_{E}\triangleq\Bigg{\{}x\in\mathcal{C}~{}\Bigg{|}\begin{array}[]{ll}\overbrace{h(\phi_{\rm b}(\tau,x))}^{\triangleq\bar{h}_{\tau}(x)}\geq 0,~{}\forall\tau\in[0,T],\\ \underbrace{h_{\rm b}(\phi_{\rm b}(T,x))}_{\triangleq\bar{h}_{\rm b}(x)}\geq 0\end{array}\Bigg{\}}.\end{split}$ (11)
@@
Note that enforcing the first constraints in (12) is not computationally tractable, as it must hold for all ${\tau\in[0,T]}$. The constraint can be discretized to a finite collection of constraints and then can be directly used for controller synthesis via the following quadratic program (BCBF-QP):
$\begin{array}[]{lll}{\mathbf{{k}^{*}}(x)=\ }\displaystyle\operatorname*{arg\,min}_{{u}\in U}\ \ {\|{u}-\mathbf{k_{d}}(x)\|^{2}}\\[8.53581pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}\textrm{s.t.}\ \ \ L_{{f}}\bar{h}_{\tau_{i}}({x})\!+\!L_{{g}}\bar{h}_{\tau_{i}}({x}){u}\!\geq\!-\alpha(h_{\tau_{i}}({x}))\\[5.69054pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}L_{{f}}\bar{h}_{\rm b}({x})\!+\!L_{{g}}\bar{h}_{\rm b}({x}){u}\!\geq\!-\alpha_{\rm b}(h_{\rm b}({x})),\end{array}$ (16)
for all ${\tau_{i}=iT/N_{\tau},~{}i=0,1,\ldots,N_{\tau}}$, where ${N_{\tau}\in\mathbb{N}}$ is the number of constraints, and ${\mathbf{k_{d}}(x):\mathbb{R}^{n}\to U}$ is a desired controller. In the upcoming experiments, the desired controller is given by the human operator’s vehicle velocity commands.
@@ II-C Deep Neural Networks
owhere $\upsilon_{\jmath}$ is the output of the $\jmath$-th layer of the DNN, and ${\upsilon_{0}=\mathbf{\bar{x}}}$ is the input to the DNN and ${\upsilon_{N_{L}}=\mathbf{\bar{y}}}$ is the output of the DNN, respectively. Each hidden layer has ${n_{\jmath}\in\mathbb{N}}$ hidden neurons. ${\mathbf{\bar{W}}_{\jmath}\in\mathbb{R}^{n_{\jmath}\times n_{\jmath-1}}}$ and ${b_{\jmath}\in\mathbb{R}^{n_{\jmath}}}$ are weight matrices and bias vectors for the $\jmath^{{th}}$ layer, respectively. ${\pi_{\jmath}\triangleq\left[\varrho_{\jmath},\cdots,\varrho_{\jmath}\right]}$ is the concentrated activation functions of the $\jmath^{{th}}$ layer wherein ${\varrho_{\jmath}:\mathbb{R}\rightarrow\mathbb{R}}$ is the activation function, such as sigmoid or ReLU:
@@
III Framework
@@
  • ${\kappa\!:\!\mathbb{R}_{\geq 0}\to\mathbb{Z}}$ is a function that identifies the index of the chosen backup strategy as a function of time.
  • ${\gamma_{t,\kappa(t)}\!\in\!\mathbb{R}^{k}}$ is a vector of ${k\!\in\!\mathbb{N}}$ features–robot state, environment, driver input, and safety data–collected while the BCBF framework utilized controller ${\mathbf{k_{b\kappa(t)}}}$.
  • We maintain an ${H\!+\!1}$-length data history ${\vec{\Gamma}(t,H,\kappa)\!=\!\left(\gamma_{t,\kappa(t)},\gamma_{t-1,\kappa(t-1)},\ldots,\gamma_{t-H,\kappa(t-H)}\right)}$. See that the chosen backup controller may change over the length of the history; it may also remain the same (for instance, $\kappa(t)$ may or may not equal $\kappa(t-1)$). Additional details on $\vec{\Gamma}(t,H,\kappa)$ are found in the following discussion.
  • $R:\mathbb{R}^{k\times(H+1)}\to\mathbb{R}^{m_{k}}$ is a reward function that maps input history $\vec{\Gamma}(t,H,\kappa)$ to a list of $m_{k}$ rewards with values ranging from 0 to 1.
@@
where ${p\!=\![{x}_{I}~{}{y}_{I}]^{\top}}$ is the vehicle’s planar position w.r.t. the inertial frame ${I}$, $\theta$ is vehicle’s yaw angle, ${-v_{\rm max}\!\leq\!v\!\leq\!v_{\rm max}}$ is its linear velocity and ${-\omega_{\rm max}\!\leq\!\omega\leq\!\omega_{\rm max}}$ is the angular velocity. While we describe our intention estimator in the context of the unicycle model, we believe that our system can generalize to more complex dynamical models. This is the subject of future experimentation as detailed in Section VI.
III-A Intention Estimation
@@
We employ an intention estimation framework to learn the reward function, $R$. Learning is useful in this context since constructing a reward function based on multi-modal input in high dimensions may be challenging. The feature vector at a single timestep, ${\gamma_{t,\kappa(t)}}$, consists of the robot position ${[x_{I},y_{I},z_{I},\theta]\in\mathbb{R}^{4}}$, robot velocities ${[\dot{x}_{I},\dot{y}_{I},\dot{z}_{I},\dot{\theta}]\in\mathbb{R}^{4}}$, driver velocity commands ${u\in U}$, and safety information from evaluating ${h:\mathbb{R}^{3}\to\mathbb{R}}$ at $x$ and at ${x_{g}\in\mathbb{R}^{3}}$, which is an intermediate goal determined by forward integrating the driver’s desired $u$ over a short horizon. These quantities are measured and calculated while the BCBF safety filter wtih backup controller $\mathbf{k_{b\kappa(t)}}$ is influencing the robot trajectory. Hence, $\kappa(t)$ is designated to the feature vector $\gamma_{t,\kappa(t)}$ to indicate this dependence. Note that $\gamma_{t,\kappa(t)}$ can contain many more features than detailed here in other implementations, especially if they relate to the safety of the robot. However, the aforementioned features generalize to any robot using our implementation. We wish to learn $R$ by mapping history $\vec{\Gamma}(t,H,\kappa)$ to rewards corresponding to each potential backup controller $\mathbf{k_{b\kappa(t+1)}}$ at next timestep $t+1$. This process allows us to derive switching laws for the next backup controller as a function of time. For instance, we may choose the backup controller with the largest reward at time $t+1$ where $\kappa(t+1)=\operatorname*{arg\,max}(R(\vec{\Gamma}(t,H,\kappa))$. This recursive formulation necessitates that we initialize $\kappa(t_{0})$ such that the robot begins in the safe control invariant set $\mathcal{C_{E}}$ corresponding to backup controller $\mathbf{k_{b\kappa(t_{0})}}$
@@
We use a deep LSTM network with a DNN decoder to construct $R$ from history $\vec{\gamma}(t,H,\kappa)$. Deep LSTMs can learn complex temporal relationships by using multiple layers and cyclic connections to understand sequential data, and the DNN decoder maps the output of the LSTM to rewards for each backup controller through several hidden layers. Both the LSTM and DNN utilize dropout for regularization, and the DNN uses ReLU activation functions in the hidden layers. Furthermore, we use the sigmoidal activation function in the DNN final layer. Thus, the outputted rewards for each backup controller, $\mathbf{k_{bi}}$, at time $t$ (where $i$ is the index of the output of $R$) can be interpreted very loosely as a likelihood that $\mathbf{k_{bi}}$ is the desired backup controller choice at time $t$. For example, if our framework is highly certain that the human operator believes $\mathbf{k_{b0}}$ is the correct choice of backup controller at time $t$, then the first component of the output of ${R(\vec{\gamma}(t,H,\kappa))}$ is expected to be close to 1.
@@
Our strategy and architecture enables easy reward engineering for the training dataset. We construct a multi-class dataset by gathering data of the robot driving in suitable environments under the correct backup controllers. During data collection the robot should be operating with a BCBF safety filter as we plan to evaluate the network in safety-critical contexts. During training, the switches between backup controllers are labeled in the dataset by the driver using configurable buttons on the robot ground station. Labeling is completed by assigning $1$ to the driver’s choice of backup controller and 00 to all others for that instance. Since we utilize the sigmoidal activation function in this multi-class classification learning task, we use the softmax loss function to train the network. Indeed, this training procedure makes an implicit assumption that there exists a single, correct choice of backup controller at any given time. While this may not be true in certain situations, we guarantee safe switching as discussed in Section III-B. Thus, any safe switch between backup controllers, does not void safety guarantees.
To improve training, we utilize the Adam optimizer [24] with a stepped learning rate scheduler. In order to compensate for the time it takes to solve the BCBF-QP, we shift the labels for the desired backup controller backwards in time. This allows the predicted backup strategy to be preemptively passed into the BCBF-QP, allowing time for the BCBF-QP to have computed safe control outputs for the new controller when they are expected.
We present the model parameters used in our final implementation and training iteration shown in Fig. 2. We used a 2 layer LSTM with 100 hidden units with the 12 total input features detailed earlier. Our DNN is composed of two hidden layers of sizes 50 and 25 neurons, respectively. We use a dropout value of 0.2 for the hidden layers of the DNN and a dropout of 0.1 for the LSTM. Accuracy is calculated by comparing the maximum reward output from the LSTM-DNN architecture and determining if it corresponds to the correct choice of backup controller for the respective point in the training set. We also use a validation (test) dataset to observe out-of-sample performance of the network during training. The network achieves 97% accuracy by the 30th epoch. We achieved this accuracy on a dataset of 19000 datapoints collected on hardware. The sequence length for the training samples of the model was chosen to be 15 timesteps, which corresponds to roughly 0.75 seconds of data. We found that this time-range produced the most accurate results.
III-B Backup Control Barrier Function Modification
Safety must be preserved when switching between backup controllers. Consider a switch from backup controller $\mathbf{k_{bi}}$ to $\mathbf{k_{bj}}$. One way to ensure safety during this crossover is to evaluate constraints (12) for the new controller $\mathbf{k_{bj}}$. Since we were previously using backup strategy $\mathbf{k_{bi}}$, we are guaranteed to be in the $T$-time reachable set of the first controller. Thus, validating that the BCBF-QP constraints hold for $\mathbf{k_{bj}}$, practically requires that both backup controllers’ $T$-time reachable sets intersect, and allows the BCBF-QP to remain solvable before and after the transition. As observed in our implementation ( IV), this resulted in relatively smooth switching between backup strategies.
III-C Human Interface